# BED-LLM: Intelligent Information Gathering with LLMs and Bayesian Experimental Design

### v1

@@ 7 Conclusion
In this work, we have shown how to effectively apply the framework of sequential Bayesian experimental design (BED) to the problem of interactive information gathering with LLMs. Specifically, we have introduced BED-LLM, which provides a specific, information–theoretic, sequential BED approach that makes a variety of carefully justified design choices in the joint-model factorization, belief updating, and candidate question generation. Particularly central to BED-LLM is the prior–likelihood pairing with filtering of hypotheses for consistency with the history. BED-LLM is notably the first work that uses both this prior–likelihood pairing without resorting to using marginal predictive entropy in lieu of the expected information gain (EIG). Together, these innovations lead to substantial performance improvements compared to previous approaches. The results thus confirm that principled EIG–driven strategies can yield substantial gains for interactive, multi–turn, information gathering problems.
Appendix A Additional discussions on design choices
A.1 An alternative view on the faithfulness of conditional distributions
@@
Another way of viewing the distinction between the prior–likelihood and data–estimation constructions is in which of the EIG forms, Equation˜2 or Equation˜3, we center our reasoning. For a given joint model, the two are, of course, mathematically equivalent. However, they give us different ways of thinking about what it means to maximize the EIG: reducing entropy in $\theta$ from seeing $y$, or reducing entropy in $y$ from seeing $\theta$. This, in turn, gives us a way to reason about how appropriate our joint model is. When we choose to use one of $p\left(y_{t+1};\left[\theta,x_{t+1}\right]\right)$ or $p\left(\theta;[h_{t},x_{t+1},y_{t+1}]\right)$, we are centering our reasoning around the entropy of this quantity making sense, while allowing the other entropy in the other form to be implicitly defined from the resulting joint distribution; because the two forms are equivalent, we know that if our explicit form is suitable/unsuitable, the implicit form will be as well. If, for example, we directly fix the form of $p\left(\theta;[h_{t},x_{t+1},y_{t+1}]\right)$ using our LLM’s predictive distribution, we are also directly relying on its expected entropy being a meaningful measure of design quality. If $\theta$ is high–dimensional and predominantly free–form, the resulting entropy produced by the LLM is unlikely to be meaningful and using the data–estimation pairing is unlikely to produce an effective strategy. However, if $y$ is instead quite constrained, the LLM can produce a meaningful entropy over it, and choosing a model based on the prior–likelihood pairing is likely to implicitly define a meaningful distribution, and thus entropy, on $\theta$. Conversely, if $\theta$ is constrained and $y$ is free form, the opposite will hold instead.
A.2 Choice of $\theta$
@@
An important corollary of this reasoning is that it can be important to be careful in our choice of exactly what we take $\theta$ to be, especially if we are using the data–estimation formulation. In particular, it is essential for entropy in the space of $\theta$ to form a meaningful notion of uncertainty, even if this entropy is not being measured through the LLM’s predictive distribution of $\theta$ directly. Thus, while $\theta$ inherently represents what we are trying to learn about and should always be set up as such, if there is flexibility in how exactly we formulate it, we should be careful to choose a form that yields an appropriate uncertainty measure. For example, if the LLM is trying to clarify what code a user wishes it to generate, we could either choose $\theta$ to be the code itself or, following (Neiswanger et al,, 2021; Bickford Smith et al,, 2023), the output the code produces. Here the entropy over code outputs induced by our distribution on code is likely to be a much better measure of uncertainty than the entropy of the raw code itself, given that there are multiple ways one can code the same operation.
A.3 Expected uncertainty reduction in the sequential setting
@@
In Section˜4.4, we looked at whether the prior–likelihood pairing or the data–estimation pairing is better aligned with how we update our beliefs in practice. The picture here can get a somewhat more complicated once we move into the sequential BED setting. Here, our ultimate aim is actually to minimize $\mathrm{H}[p(\theta;h_{T})]$ at some final future horizon $T$. Now, we only care about intermediary belief states $p(\theta;h_{t})$ through their aid in future decision making toward the goal of minimizing the final entropy. Thus, even if we are working with in-context updates, it might be the case that $p(\theta;h_{t})$ only starts to produce a meaningful entropy once we have seen enough data to sufficiently narrow down the possibilities on $\theta$. The optimal behavior in such settings would be to learn a policy that directly targets this final belief state instead of sequentially targeting the incremental EIGs. However, this will typically not be computationally feasible in practice and we instead need to resort to a myopic decision-making strategy. It might thus still be better to use the prior–likelihood formulation in such myopic decision making settings, even if we are sequentially updating our beliefs on $\theta$ through in-context updates, if this allows us to better guide the sequential decisions towards our final objective. The coherence of Bayesian updating means that the converse is unlikely to be true, so this provides further evidence towards using the prior–likelihood formulation.
Appendix B Generating candidate hypotheses for BED-LLM
A fundamental challenge for BED-LLM and the same method with the Entropy baseline is generating a sufficiently diverse set of candidate hypotheses from the LLM prior. As illustrated in Figure˜8, the raw distribution $p_{\mathrm{LLM}}(\theta)$ is highly over–collapsed, often concentrating mass on only a few high–likelihood hypotheses. Thus, it is not practical to directly use the LLM’s distribution as a prior $p(\theta)$. Instead, we jointly sample candidate $\theta$ and assume a uniform distribution over them. We can view this as sampling thetas from a mixture distribution.
Instead, we employ a joint sampling strategy, in which the LLM is prompted to generate a list of $N$ hypotheses in a single rollout, which corresponds to drawing from the autoregressive list distribution
$p_{\mathrm{LLM}}(\theta^{(1)}_{t},\ldots,\theta^{(N)}_{t};h_{t})\;=\;\prod_{n=1}^{N}p_{\mathrm{LLM}}(\theta_{t}^{(n)};[\theta_{t}^{(1:n-1)},h_{t}])$
Each such joint sample is then filtered for history consistency using $p_{\mathrm{LLM}}(h_{t})$. This may be repeated. The surviving hypotheses from the current rollout are then merged with those retained from previous steps to form a cumulative candidate set $\Theta$. While one could in principle reweight candidates using importance sampling, in practice we choose to not rely on the model’s internal probabilities. Instead approximate the prior as a uniform distribution over this union
$p_{f}(\theta;h_{t})\approx\frac{1}{\Theta}\sum_{\theta\in\Theta}\delta_{\theta}.$
To further improve coverage and diversity, we incorporate several refinements in the generation of $p_{f}(\theta;h_{t})$. Due to the differences between different LLMs, the ways this is encouraged may differ, but we find in general that the following are effective:
  • Repeat sampling and filtering: The sampling, filtering, and merging steps may be repeated to encourage diversity and ensure filtering correctness.
  • Prompts: Prompts are designed to elicit stratified hypotheses by encouraging the LLM to consider different semantic features (e.g. age groups, genres, or categories) and implicitly diversify across them.
  • Context ordering: During generation, we reverse the order of the question–answer pairs in $h_{t}$ to place the most recent question at the top of the context window (while retaining earlier exchanges), ensuring that specific constraints are prioritized and mitigating context drift.
Finally, we note that different LLMs respond differently to strategies aiming to increase diversity: some benefit more from a higher temperature while others benefit from more repetitions of the sampling–filtering cycle.
Appendix C Data–estimation method
C.1 EIG Estimation
Suppose our model is given by $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p\left(y_{t+1};[h_{t},x_{t+1}]\right)p\left(\theta;[h_{t},x_{t+1},y_{t+1}]\right)$}. Here, it will clearly be beneficial to directly use Equation˜2 for estimating the EIG, as here we directly have access to all the required terms, other than $p(\theta)$ which can be simply ignored as it is not a function of $x$ so does not affect optimization of the question. If the possible values for $y$ are enumerable and we can evaluate $p\left(y_{t+1};[h_{t},x_{t+1}]\right)$ in closed–form, we can directly calculate the exact EIG (up to a constant) without requiring any estimation at all:
$\text{EIG}_{\theta}(x)-\mathrm{Const}=-\sum_{y}p\left(y_{t+1};[h_{t},x_{t+1}]\right)\mathrm{H}\left[p\left(\theta;[h_{t},x_{t+1},y_{t+1}]\right)\right],$ (8)
where the entropy $\mathrm{H}\left[p\left(\theta;[h_{t},x_{t+1},y_{t+1}]\right)\right]$ can be evaluated directly from the logits of the LLM, or if these are not available, estimated by sampling. If we cannot enumerate $y$ or evaluate $p\left(y_{t+1};[h_{t},x_{t+1}]\right)$, we can simply instead resort to Monte Carlo and use the estimator:
$\text{EIG}_{\theta}(x)-\mathrm{Const}\approx-\frac{1}{N}\sum_{n=1}^{N}\mathrm{H}\left[p(\theta;[c,x,y_{n}])\right]\quad\mathrm{where}\quad y_{n}\sim p\left(y_{t+1};[h_{t},x_{t+1}]\right).$ (9)
@@ C.1 EIG Estimation
1:History $h_{t}$; candidate questions $\mathcal{X}_{\text{cand}}^{t}$; answer sets $\{\mathcal{Y}(x)\}$
@@
4:  Obtain predictive answer distribution $p_{\text{LLM}}(y;x,h_{t})$ for all $y\in\mathcal{Y}(x)$
@@
6:   Construct $\Theta_{t,y}\sim p_{f}(\theta;h_{t}\cup(x,y))$
7:   Compute entropy $H_{y}\leftarrow\mathrm{H}\!\left[p_{\mathrm{LLM}}(\theta;h_{t},x,y)\right]$
8:  end for
9:  Compute $\text{EIG}(x)\leftarrow-\sum_{y\in\mathcal{Y}(x)}p_{\text{LLM}}(y;x,h_{t})\;H_{y}$
10:end for
11:Select $x_{t}^{\star}\leftarrow\arg\max_{x\in\mathcal{X}_{\text{cand}}^{t}}\text{EIG}(x)$
12:return $x_{t}^{\star}$
C.2 Generating candidate hypotheses
To generate candidate values of $\theta$ for the data–estimation method, we used the prompt in Figure˜7. Figure˜8 shows an example of the distribution of samples obtained following two rounds in the 20–questions game. Note that the samples are highly concentrated on just a handful of answers. This lack of diversity shows that the model’s belief distribution is far more concentrated relative to the variability over valid hypotheses in the ground–truth task distribution, which negatively impacts the performance of the data–estimation method.
Return only the full name of one randomly selected famous person (living or deceased) consistent with the questions and answers above.

### v2

@@ 7 Conclusion
In this work, we have shown how to effectively apply the framework of sequential Bayesian experimental design (BED) to the problem of interactive information gathering with LLMs. Specifically, we have introduced BED-LLM, which provides a specific, information–theoretic, sequential BED approach that makes a variety of carefully justified design choices in the joint-model factorization, belief updating, and EIG estimation. Particularly central to BED-LLM is the prior–likelihood pairing with filtering of hypotheses for consistency with the history. BED-LLM is notably the first work that uses both this prior–likelihood pairing without making a deterministic likelihood assumption that causes the EIG to simply to just marginal predictive entropy. Together, these innovations lead to substantial performance improvements compared to previous approaches. The results thus confirm that principled EIG–driven strategies can yield substantial gains for interactive, multi-turn, information gathering problems.
Appendix Contents
Appendix A Algorithmic considerations
A.1 Updating the likelihood
The success of BED-LLM hinges on our ability to update our joint distribution. As mentioned in §3, we choose not to update the likelihood model as more data is gathered, that is, our likelihood in the sequential setting will be $p_{\mathrm{LLM}}\left(y_{t};\left[\theta,x_{t}\right]\right)$ instead of $p_{\mathrm{LLM}}(y_{t+1};[h_{t-1},\theta,x_{t+1}])$. The main rationale of this choice is that for many problems our beliefs on $\theta$ capture all the required information to predict $y|x$, hence including the history is adding unnecessary context that could influence the LLM’s behavior in undesirable ways. However, it is important to note that $p_{\mathrm{LLM}}(y_{t+1};[h_{t-1},\theta,x_{t+1}])$ should be used instead for problems where $\theta$ will not capture all information from previous data, e.g. if $\theta$ is a binary value corresponding to whether we reject a null hypothesis, or is the answer to a particular other question of interest.
A.2 Estimating EIG for each question
Question 1 Question:
Which ice cream flavor feels like the best match for this user?
Predictive Entropy: Very High
EIG: 0
Question 2 Question:
Which film genre does the user most prefer?
Choose one option: A. Action B. Sci-Fi C. Comedy D. Horror
Predictive Entropy: High
EIG: High
A.2.1 Predictive entropy is not a good approximation for EIG
As discussed in §3.2, previous information-based query selection mechanisms have assumed that responses are deterministic given $\theta$ and $x$. This implies that the expected entropy of the likelihood, $\mathbb{E}_{p(\theta;h_{t-1})}[\mathrm{H}[p(y_{t+1}|\theta;x_{t+1},h_{t-1})]]$, is constant over designs, meaning that maximizing EIG is equivalent to maximizing the marginal predictive entropy, $\mathrm{H}[\mathbb{E}_{p(\theta;h_{t-1})}[p(y_{t}|\theta;x_{t},h_{t-1})]]$.
In practice, the expected likelihood entropy can and will vary across designs. This variability in the expected likelihood entropy can be crucial in selecting good designs. Here, we walk through a concrete example where predictive entropy might differ significantly from EIG.
Figure˜3 shows two candidate questions that could be asked to elicit film preference. Question 1 has high predictive entropy: in a randomly selected group of people, we would expect high variation in ice cream preference (regardless of the individual’s film preferences). However, since ice cream preference is unrelated to film preference, the answer would not help us narrow down our hypothesis space, and the EIG is zero.
This is also supported by evidence in our experiments (§6). Both the Split baseline, and the Entropy ablation, assume a deterministic likelihood; in particular, the Entropy ablation uses the same estimator of the predictive entropy as BED-LLM. In both cases, we see the performance significantly degrades relative to using the full EIG. Further, omitting the expected likelihood entropy term provides no meaningful computational saving—the same LLM evaluations are used for the top and bottom lines of Equation˜3, hence doing the full estimate of the EIG requires no additional LLM calls to be made.
A.2.2 EIG estimator
One might be tempted to replace $\hat{p}(y_{t+1};[h_{t-1},x_{t+1}])$ with $p_{\mathrm{LLM}}(y_{t+1};[h_{t-1},x_{t+1}])$ in the EIG estimation in Equation˜3, as the two essentially offer alternative predictive distributions for the outcome. We also advise against this though, noting that it again provides no meaningful computational benefits (unless one also assumes a deterministic likelihood, but this would then mean we no longer consider $\theta$ at all). A key reason for avoiding this substitution is that it would mean we are no longer estimating a true EIG: the inconsistency between the likelihood and the marginal data distribution means there is no longer a joint model where we are minimising our expected uncertainty in $\theta$. We also find that the LLM process of sampling $\theta$ from $p_{f}(\theta;h_{t-1})$ followed by $y$ from $p_{\mathrm{LLM}}\left(y_{t};\left[\theta,x_{t}\right]\right)$ tends to give a better uncertainty over responses than sampling $y$ directly from $p_{\mathrm{LLM}}(y_{t+1};[h_{t-1},x_{t+1}])$.
A.3 Prior construction and belief updating
In Section˜3.3, we argued that naive in-context updating is not sufficient for updating our beliefs: We fail to fully incorporate the information from the history $h_{t}$, and we often have overconfident distributions. The shortfalls of in-context learning in such settings have also previously be noted by, for example, (Liu et al., 2024b; Zhang et al., 2025; 2024). We posit two reasons why they likely struggle in such settings. First, the information from the different examples in the history are generally highly distinct in these information-gathering settings (indeed, this is part of our aim in adaptively design informative questions), making it harder for the LLM to appropriately reconcile all the provided information than in many other uses of in–context learning. Second, $\theta$ will often represent a user–specific variable that cannot easily be predicted from any data other than the user’s responses to questions: it has been argued that much of the success of in–context learning in LLMs is down to improving problem specification and linking the requested task to data it has seen in its training, rather than truly “learning” from the provided examples (Min et al., 2022; Kossen et al., 2024), but the history in our setting is rarely helpful for this due to its user–specific nature.
Appendix B Additional discussions on design choices
B.1 An alternative view on the faithfulness of conditional distributions
@@
Another way of viewing the distinction between the prior–likelihood and data–estimation constructions is in which of the EIG forms, Equation˜1 or Equation˜2, we center our reasoning. For a given joint model, the two are, of course, mathematically equivalent. However, they give us different ways of thinking about what it means to maximize the EIG: reducing entropy in $\theta$ from seeing $y$, or reducing entropy in $y$ from seeing $\theta$. This, in turn, gives us a way to reason about how appropriate our joint model is. When we choose to use one of $p\left(y_{t};\left[\theta,x_{t}\right]\right)$ or $p\left(\theta;[h_{t-1},x_{t},y_{t}]\right)$, we are centering our reasoning around the entropy of this quantity making sense, while allowing the other entropy in the other form to be implicitly defined from the resulting joint distribution; because the two forms are equivalent, we know that if our explicit form is suitable/unsuitable, the implicit form will be as well. If, for example, we directly fix the form of $p\left(\theta;[h_{t-1},x_{t},y_{t}]\right)$ using our LLM’s predictive distribution, we are also directly relying on its expected entropy being a meaningful measure of design quality. If $\theta$ is high–dimensional and predominantly free–form, the resulting entropy produced by the LLM is unlikely to be meaningful and using the data–estimation pairing is unlikely to produce an effective strategy. However, if $y$ is instead quite constrained, the LLM can produce a meaningful entropy over it, and choosing a model based on the prior–likelihood pairing is likely to implicitly define a meaningful distribution, and thus entropy, on $\theta$. Conversely, if $\theta$ is constrained and $y$ is free form, the opposite will hold instead.
B.2 Choice of $\theta$
@@
An important corollary of this reasoning is that it can be important to be careful in our choice of exactly what we take $\theta$ to be, especially if we are using the data–estimation formulation. In particular, it is essential for entropy in the space of $\theta$ to form a meaningful notion of uncertainty, even if this entropy is not being measured through the LLM’s predictive distribution of $\theta$ directly. Thus, while $\theta$ inherently represents what we are trying to learn about and should always be set up as such, if there is flexibility in how exactly we formulate it, we should be careful to choose a form that yields an appropriate uncertainty measure. For example, if the LLM is trying to clarify what code a user wishes it to generate, we could either choose $\theta$ to be the code itself or, following (Neiswanger et al., 2021; Bickford Smith et al., 2023), the output the code produces. Here the entropy over code outputs induced by our distribution on code is likely to be a much better measure of uncertainty than the entropy of the raw code itself, given that there are multiple ways one can code the same operation.
B.3 Alignment between EIG and belief updating procedure
Our ultimate goal is to minimize uncertainty in $\theta$, as measured by its entropy. With this in mind, we can use the expected uncertainty reduction framework of Bickford Smith et al. (2025) to provide insights into how well our EIG formulation and belief updating procedures align.
To simplify discussions, for now we consider the setting where we choose a single question $x$ and obtain a response $y$. Following Bickford Smith et al. (2025), we can think of the “true” optimal design as selecting
Note here that true optimal design has no direct dependency on our current beliefs about $\theta$; it only depends on $p_{\mathrm{true}}(y;x)$ and the hypothetical beliefs we produce for given observed data, $p(\theta;x,y)$. Thus, we can now see that our choice of joint model corresponds to different choices for approximating these quantities. Assuming that the LLM distribution is used directly for the conditional as per Section˜4, we thus have that:
The former of these considerations is difficult to control for as we simply do not know the true response distribution and it is hard to say which approach will thus estimate it best (though we can refer to the discussion in Section˜4 to determine which best matches our beliefs about the true response distribution). However, we do know upfront how we plan to derive our belief distribution on $\theta$ in practice, so we can use this to guide which joint model we formulate our EIG from. Namely, we observe that: a) using the prior–likelihood EIG pairing equates to assuming we will make a Bayesian update to our beliefs on $\theta$ using the likelihood $p_{\mathrm{LLM}}(y;[\theta,x])$; b) using the data–estimation EIG pairing equates to assuming we will make an in-context update to our beliefs on $\theta$, as we are treating $p(\theta;x,y)$ as $p_{\mathrm{LLM}}(\theta;[x,y])$.
For the settings we consider, we expect $p_{f}(\theta;[x,y])$ to generally be better approximated by a Bayesian update than an in–context update, aligning with our decision to use the prior–likelihood formulation. The reasons for this are that a) the filtering often removes a large proportion of the generated samples, especially at later experiment turns, with $p_{\mathrm{LLM}}(\theta;h_{t-1})$ not fully incorporating information from the history; b) the maintaining of the set of one consistent hypotheses from one turn to the next encourages a more Bayesian behavior, with samples persisting unless contradicted by a new likelihood term; and c) the typical premature overconfidence of $p_{\mathrm{LLM}}(\theta;h_{t-1})$ to a small number of hypotheses means it is typically unrepresentative of our beliefs.
@@ B.3 Alignment between EIG and belief updating procedure
The picture here can get a somewhat more complicated once we move into the sequential BED setting. Here, our ultimate aim is actually to minimize $\mathrm{H}[p(\theta;h_{T})]$ at some final future horizon $T$. Now, we only care about intermediary belief states $p(\theta;h_{t})$ through their aid in future decision making toward the goal of minimizing the final entropy. Thus, even if we are working with in-context updates, it might be the case that $p(\theta;h_{t})$ only starts to produce a meaningful entropy once we have seen enough data to sufficiently narrow down the possibilities on $\theta$. The optimal behavior in such settings would be to learn a policy that directly targets this final belief state instead of sequentially targeting the incremental EIGs. However, this will typically not be computationally feasible in practice and we instead need to resort to a myopic decision-making strategy. It might thus still be better to use the prior–likelihood formulation in such myopic decision making settings, even if we are sequentially updating our beliefs on $\theta$ through in-context updates, if this allows us to better guide the sequential decisions towards our final objective. The coherence of Bayesian updating means that the converse is unlikely to be true, so this provides further evidence towards using the prior–likelihood formulation.
Appendix C Extended related work
Information-based question answering with LLMs
Several recent works have (explicitly or implicitly) looked at information gathering with LLMs. Most of these can be framed in a BED setting, with a deterministic likelihood (Piriyakulkij et al., 2023; Hu et al., 2024; Kobalczyk et al., 2025; Cooper et al., 2025), and can be seen as variants of our Split baseline. Piriyakulkij et al. (2023) use a deterministic 0/1 answer likelihood $p(a|x,q)$ via the LLM to prune items from a pre-enumerated finite set given a candidate question $q$. The question is selected by minimizing expected posterior entropy. They model user preferences with a binary ground truth, which would not be applicable in preference-elicitation scenarios with nebulous user profiles. Similarly,  Hu et al. (2024) use a deterministic likelihood to minimize entropy over a finite set $\Omega$ in a closed-world setting. Kobalczyk et al. (2025) target ambiguous task specifications in open-ended generation tasks by sampling a set of hypotheses (placing a uniform prior over them) and viewing each question as a deterministic partition over those samples, looking for questions that split the samples roughly evenly. Cooper et al. (2025) compute posterior entropy over a working set of top-k hypotheses (without filtering) through heuristic pruning.
Wang et al. (2025) avoid the pitfall of deterministic likelihoods. They use a data–estimation framework to estimate EIG, focusing on scenarios where the target can be expressed as a predefined series of multiple-choice questions. Their approach relies on meta-training a predictive language model on historic question/answer pairs, and so is not directly comparable with BED-LLM which requires no additional training or data. Chan et al. (2025) do not model likelihoods or posterior beliefs, instead they rely on the expected size of conformal prediction sets as a surrogate uncertainty metric. This requires the use of an additional calibration dataset, and is confined to closed-world settings with a finite label set and pre-defined queries.
Post-training LLMs for improved information gathering
Rather than augmenting a frozen LLM with the ability to estimate utility functions, some works have instead aimed to post-train an LLM to improve its ability to ask questions (Zhang et al., 2024; Wu et al., 2025; Andukuri et al., 2024). Most do not explicitly consider informativeness of questions: Zhang et al. (2024) and Wu et al. (2025) use reinforcement learning techniques to reward generations that quickly lead to the correct answer, and Andukuri et al. (2024) builds on Li et al. (2025b) by fine-tuning on successful traces. Mazzaccara et al. (2024) do indirectly incorporate uncertainty, also using a deterministic likelihood: they use predictive entropy to identify informative questions, and then either fine-tune on the highest-entropy question, or perform DPO comparing the highest-entropy question with a lower-entropy question. We do not address fine-tuning in this work, focusing instead on exploring the correct way to formulate BED using LLMs.
Combining LLMs with parametric models
As discussed in Section˜4, a key challenge in adapting BED to the LLM setting is in aligning the expected information gain with the actual uncertainties extracted from the LLM after updating. Handa et al. (2024) take a different approach to this problem by using the LLM to generate features for an external conventional Bayesian joint model (in their case, a linear Bradley–Terry model), rather than deriving their joint model more directly from the LLM itself. This can be a good choice when the problem is well-bounded and we already have a well-specified Bayesian model form for the problem at hand; however, this may be challenging in arbitrarily large and complex hypothesis spaces. In particular, their specific method is not applicable more widely beyond the preference learning context they consider.
BED
It has been noted that the traditional sequential BED approach can sometimes be suboptimal in practice, as it only optimizes the EIG of the next observation, without planning ahead for the fact that design decisions taken at a given step can also influence the achievable EIGs from future steps (Foster, 2021). A variety of policy–based BED approaches have subsequently been proposed to address this (Foster et al., 2021; Ivanova et al., 2021; Blau et al., 2022; Huan & Marzouk, 2016; Hedman et al., 2025), while also removing the need to make model updates and conduct optimizations during the experiment itself. Our findings are complementary: by providing more faithful model factorizations, belief updates, and EIG estimators in the LLM setting, BED-LLM could supply stronger building blocks for policy-based methods, reducing variance, enhancing effectiveness, and improving the sample efficiency of policy training.
@@ Appendix C Extended related work
1:History $h_{t-1}$; candidate questions $\mathcal{X}_{\text{cand}}^{t}$; answer sets $\{\mathcal{Y}(x)\}$
@@
4:  Obtain predictive answer distribution $p_{\text{LLM}}(y;x,h_{t-1})$ for all $y\in\mathcal{Y}(x)$
@@
6:   Compute entropy $H_{y}\leftarrow\mathrm{H}\!\left[p_{f}(\theta;h_{t-1}\cup(x,y))\right]$
7:  end for
8:  Compute $\text{EIG}(x)\leftarrow-\sum_{y\in\mathcal{Y}(x)}p_{\text{LLM}}(y;x,h_{t-1})\;H_{y}$
9:end for
10:Select $x_{t}^{\star}\leftarrow\arg\max_{x\in\mathcal{X}_{\text{cand}}^{t}}\text{EIG}(x)$
11:return $x_{t}^{\star}$
Appendix D Data–estimation method
D.1 EIG Estimation
Suppose our model is given by $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p\left(y_{t};[h_{t-1},x_{t}]\right)p\left(\theta;[h_{t-1},x_{t},y_{t}]\right)$}. Here, it will clearly be beneficial to directly use Equation˜1 for estimating the EIG, as here we directly have access to all the required terms, other than $p(\theta)$ which can be simply ignored as it is not a function of $x$ so does not affect optimization of the question. If the possible values for $y$ are enumerable and we can evaluate $p\left(y_{t};[h_{t-1},x_{t}]\right)$ in closed–form, we can directly calculate the exact EIG (up to a constant) without requiring any estimation at all:
$\text{EIG}_{\theta}(x)-\mathrm{Const}=-\sum_{y}p\left(y_{t};[h_{t-1},x_{t}]\right)\mathrm{H}\left[p\left(\theta;[h_{t-1},x_{t},y_{t}]\right)\right],$ (5)
where the entropy $\mathrm{H}\left[p\left(\theta;[h_{t-1},x_{t},y_{t}]\right)\right]$ can be evaluated directly from the logits of the LLM, or if these are not available, estimated by sampling. If we cannot enumerate $y$ or evaluate $p\left(y_{t};[h_{t-1},x_{t}]\right)$, we can simply instead resort to Monte Carlo and use the estimator:
$\text{EIG}_{\theta}(x)-\mathrm{Const}\approx-\frac{1}{N}\sum_{n=1}^{N}\mathrm{H}\left[p(\theta;[c,x,y_{n}])\right]\quad\mathrm{where}\quad y_{n}\sim p\left(y_{t};[h_{t-1},x_{t}]\right).$ (6)
D.2 Generating candidate hypotheses
To generate candidate values of $\theta$ for the data–estimation method, we use the prompt in Section˜D.2.
Figure 4: Prompt for generating hypotheses (and evaluating their probability) for the data–estimation method. Return only the full name of one randomly selected famous person (living or deceased) consistent with the questions and answers above.